NvidiaIntely (nvidiaintely):

At a certain time of day, a 6 foot tall man has a 8 foot shadow. If a tree is 25 feet tall, what is the length of the tree’s shadow? (Round to the nearest hundredth)

NvidiaIntely (nvidiaintely):

@Kevin

OpenStudy (kevin):

How... how do you know I'm online o_O

jhonyy9 (jhonyy9):

i ve posted to you before the right way to solve it

NvidiaIntely (nvidiaintely):

Yeah but I posted again, because my PC crashed and I lost that post :(

jhonyy9 (jhonyy9):

Kevin just push the coursor on your wrote name and will write than you are online or off line

OpenStudy (kevin):

Yeah, I know xD lol

jhonyy9 (jhonyy9):

6 foot tall man --- 8 ft shadow 25 ft tall tree ----- x ft shadow ------------------------------ x = 25*8/6 = ? ft

NvidiaIntely (nvidiaintely):

Okay thnaks @jhonyy9, but how do I solve for x? ik you show it there, but how do I do that?

jhonyy9 (jhonyy9):

x = (25*8)/6 = 200/6 = ?

NvidiaIntely (nvidiaintely):

what does / mean?

NvidiaIntely (nvidiaintely):

Is it divide?

NvidiaIntely (nvidiaintely):

Or fraction?

jhonyy9 (jhonyy9):

yes

NvidiaIntely (nvidiaintely):

Okay let me try it :)

NvidiaIntely (nvidiaintely):

Is it 33.30?

jhonyy9 (jhonyy9):

33,33 this is the right answer

NvidiaIntely (nvidiaintely):

Okay. But it says round to the nearest hundredth.

jhonyy9 (jhonyy9):

rounded will be 33,35

NvidiaIntely (nvidiaintely):

Thanks Jhonyy!

jhonyy9 (jhonyy9):

yw

OpenStudy (welshfella):

6/8 = 25 / x x = 25*8 / 6 = 33.33.... to nearest hundredth its 33.33

NvidiaIntely (nvidiaintely):

yeah, thanks!

jhonyy9 (jhonyy9):

rounded why not is 33,35 ?

jhonyy9 (jhonyy9):

@mathmale

jhonyy9 (jhonyy9):

just bc. you get 33,333333 and not 33,3222222

OpenStudy (welshfella):

these are similar triangles and the corresponding sides are in the same ratio 

jhonyy9 (jhonyy9):

ok. i understand this but why not is rounded 33,35 and just 33,33 ?

OpenStudy (welshfella):